# Conformally invariant boundary arcs in double dimers

### v1

@@ Appendix B Inverse Kasteleyn matrix in Temperleyan domains
In this section, the inverse Kasteleyn matrix on Temperleyan domains will often be called the coupling function as in the original work of [15]. We detail why it also works near the boundary. We use the notation and setting of Section 3. We start by recalling some well-known facts on the coupling function and discrete holomorphic functions. The acquainted reader can jump directly to the statements of Lemma B.2 and Lemma B.3.
@@
where $G=\Delta^{-1}$ is the Green’s function. The function $G$ takes real values because the Laplacian $\Delta$ takes real values. Further, since the Laplacian has a block structure (it is non zero only from $B_{0}$ to $B_{0}$ and $B_{1}$ to $B_{1}$), $G$ has the same block structure. Hence, since $w\in W_{0}$, $w\pm\frac{\varepsilon}{2}\in B_{0}$ while $w\pm\frac{i\varepsilon}{2}\in B_{1}$, so $\tilde{K}^{-1}(b,w)$ is real for $b\in B_{0}$ and pure imaginary for $b\in B_{1}$ (justifying the terminology “real” and “imaginary” part).
Let us define $\partial^{out}B_{0}$ (the full red and blue circles on Figure 15) to be the vertices in $B_{0}(\varepsilon\mathbb{Z}^{2})\setminus B_{0}$ (where $B_{0}(\varepsilon\mathbb{Z}^{2})$ are the vertex of the dual graph of $\varepsilon\mathbb{Z}^{2}$ i.e. the faces of $\varepsilon\mathbb{Z}^{2}$) at distance $\varepsilon$ from a vertex of $B_{0}$ (see Kenyon, section 4.1) and $\partial^{out}B_{1}=\{b_{0}\}$ (full red square on Figure 15). Define $\overline{B_{0}}=B_{0}\cup\partial^{out}B_{0}$, $\overline{B_{1}}=B_{1}\cup\partial^{out}B_{1}$ and $\overline{B}=\overline{B_{0}}\cup\overline{B_{1}}$. If we extend $b\in B_{0}\to\tilde{K}^{-1}(b,w)$ by 00 on the vertices of $\partial^{out}B_{0}$, it is a harmonic function on $B_{0}$ except at two points: $\Delta\tilde{K}^{-1}(\cdot,w)_{|B_{0}}=\delta_{w+\varepsilon/2}-\delta_{w-\varepsilon/2}$. Indeed, for $b\in B_{0}$,
@@
as a function of $b\in B_{0}$. Observe that this already fully characterizes the real part: there is a unique function which satisfies (45) and is 00 on $\partial^{out}B_{0}$ (by the maximum principle). If we extend the imaginary part $b\in B_{1}\to\tilde{K}^{-1}(b,w)$ by 00 on $b_{0}$, the imaginary and real part are harmonically conjugated: $\tilde{K}\tilde{K}^{-1}=I$ implies that for $w^{\prime}\neq w$,
@@
In our case (and unlike Kenyon who considers multiply connected domains), the imaginary part is fully defined by saying that it is the harmonic conjugate of the real part and it is 00 at the removed vertex $b_{0}$: by discrete integration along paths which avoid the singularity at $w$, the values of $\tilde{K}^{-1}(\cdot,w)$ on $B_{1}$ can be uniquely recovered from the values on $B_{0}$ and the boundary conditions at $b_{0}$. We summarize all this in a lemma:
@@
  • its real part is the (unique) harmonic function on $\overline{B_{0}}$ which is 00 on $\partial^{out}B_{0}$ (Dirichlet boundary conditions) and satisfies
@@
We are now ready to give estimates of the coupling function. Let $\tilde{g}(u,v)$ be the analytic function of $v$ whose real part is the Dirichlet Green’s function $g(u,v)$ on $U$. Also define, as in Section 6.3 of [17],
@@
$\tilde{K}^{-1}_{s}(s(b),w)=\left\{\begin{array}[]{ll}-\tilde{K}^{-1}(w,b)&\text{if }b\in\overline{B_{0}}\\ \tilde{K}^{-1}(w,b)&\text{if }b\in\overline{B_{1}}\\ \end{array}\right.$ (49)
@@
We now show that the extended coupling function $b\in B^{s}\to\tilde{K}^{-1}_{s}(b,w)$ coincides with the difference of two coupling functions on $B^{s}$. Since $\mathcal{G}_{\varepsilon}^{s}$ is a Temperleyan approximation of $U\cup s(U)$ if we remove the vertex $b_{0}$, by Lemma B.1 for fixed $w\in W_{0}\cup s(W_{0})$, the coupling function $b\in B^{s}\to H^{s}(b,w)$ is uniquely defined: its real part has 00 boundary conditions on $\partial^{out}B_{0}\cup s(\partial^{out}B_{0})\cup\{b_{l},b_{r}\}\setminus\partial^{flat}B_{0}$ and satisfies $\Delta H^{s}(\cdot,w)=\delta_{w+\varepsilon/2}-\delta_{w-\varepsilon/2}$ at all inner points $b\in B^{s}_{0}$. Its imaginary part is 00 on $b_{0}$ and is harmonically conjugated with the real part except at $w$. We claim that for all $b\in B^{s}_{0}$,
@@
Observe that we did not impose any condition at $s(b_{0})$ for $H^{s}(b,w)-H^{s}(b,s(w))$, but since equality (50) holds it must be 00 a posteriori. This can be checked using symmetry arguments.
@@
If we denote by $\widetilde{g^{s}}(u,v)$ the analytic function of $v$ whose real part is the Dirichlet Green’s function $g^{s}(u,v)$ on $U\cup s(U)$, we define $F_{+}^{s}$ and $F^{s}_{-}$ similarly to the definition before Lemma B.2. If $w$ is at distance $O(\varepsilon)$ from $u$ and $b$ is at distance $O(\varepsilon)$ from $v$,
@@
where the $O(\varepsilon^{2})$ is uniform once $\delta$ is fixed. Observe now (this is the continuous equivalent of what we have just done, i.e. Schwarz reflection principle in the continuum) that $\widetilde{g^{s}}(u,v)-\widetilde{g^{s}}(s(u),v)=\tilde{g}(u,v)$ when $u,v\in U$, because the left-hand side is analytic in $v$ on $U$, its real part $g^{s}(u,v)-g^{s}(s(u),v)$ has the same pole as $g(u,v)$ in $v$ (by definition as a Green’s function) and it is 00 on the boundary of $U$. On the flat portion, it is true by symmetry: if $v$ is on the reflection axis, $g^{s}(u,v)=g^{s}(s(u),s(v))=g^{s}(s(u),v)$. On the rest of the boundary both terms of the difference vanish since $v$ is also on the boundary of $U\cup s(U)$. Hence, $F_{+}^{s}(u,v)-F_{+}^{s}(s(u),v)=F_{+}(u,v)$ and $F_{-}^{s}(u,v)-F_{-}^{s}(u,v)=F_{-}(u,v)$ which concludes the proof in the case $w\in W_{0}$; Lemma B.2 holds up to the boundary.
To deal with the case $w\in W_{1}$, the same argument works but now $\tilde{K}^{-1}(\cdot,w)_{|B_{0}}$ is the unique function on $B_{0}$ which is 00 on $\partial^{out}B_{0}$ and whose Laplacian is $-i\delta_{b=w+i\varepsilon/2}+i\delta_{w-i\varepsilon/2}$ by equation (45) applied when $w\in W_{1}$ instead of $W\in W_{0}$. The remaining changes are straightforward. ∎
@@
We use the setting of Section 3: $U^{r}$ is a simply connected open set symmetric with respect to the horizontal axis, $U=U^{r}\cap\mathbb{R}\times\mathbb{R}_{>0}$. The boundary of $U$ can be partitioned in two connected arcs: $\partial U=D\sqcup N$ where $N=\partial U\cap\mathbb{R}\times\{0\}$ is the intersection with the horizontal axis. Recall that we defined two graphs approximating $U$: the Temperleyan approximation $\mathcal{G}_{\varepsilon}^{1}$ with Dirichlet boundary conditions (and a vertex $b_{0}$ removed) and the piecewise Temperleyan approximation $\mathcal{G}_{\varepsilon}^{2}$ with two convex white corners $v_{1}^{*},v_{2}^{*}$. In both cases, we use the usual Kasteleyn phases of discrete holomorphy described in Equation (3), and denote by $K^{1},K^{2}$ the corresponding Kasteleyn matrices on $\mathcal{G}_{\varepsilon}^{1},\mathcal{G}_{\varepsilon}^{2}$. Both $\mathcal{G}_{\varepsilon}^{1}$ and $\mathcal{G}_{\varepsilon}^{2}$ have at least one dimer configuration (for example because from any loops and arcs configuration $\omega\in\Omega(\mathcal{G}_{\varepsilon})$ we can obtain two dimer covers of $\mathcal{G}_{\varepsilon}^{1}$ and $\mathcal{G}_{\varepsilon}^{2}$), hence $K^{1}$ and $K^{2}$ are invertible by [20]. Moreover, the asymptotics of $(K^{j})^{-1}$ on piecewise Temperleyan domains are computed and expressed in terms of $f_{i}^{j}(u,v)$ for $i\in\{0,1\},~{}j\in\{1,2\}$. These are the unique functions which are analytic in $u$ for fixed $v$, have a simple pole of residue $1/\pi$ at $u=v$ with no other pole on $\bar{U}$, are bounded near $v_{1}^{*},v_{2}^{*}$ (for the $j=2$ case), and have the following boundary conditions:
@@
where $\lambda=e^{i\frac{\pi}{4}}$ and $F^{\varepsilon}_{\mathbb{C},w}$ is the unique discrete holomorphic function on $\varepsilon\mathbb{Z}^{2}$ tending to 00 at infinity and with a pole $\lambda$ at $w$, which is known to be asymptotically equal to $\frac{\lambda}{2\pi(b-w)}$ when $\varepsilon\to 0$ and $|b-w|\geq\eta$ for some $\eta>0$ independent of $\varepsilon$. See Section 5 of [20] for details. The error term $\Delta$ is also present in the original statement of Lemmas B.2 and B.3, but we included it in the $o(\varepsilon)$ since we never use these lemmas for close points and for all $\eta>0$, for all $b,w$ such that $|b-w|\geq\eta$,
@@
As in Lemma B.3, when $U$ has some specific boundary conditions, the asymptotic estimates holds up to the boundary.
@@
The same argument as in the proof of Lemma B.3 can be applied: we skip most of the details. For fixed $w\in W(\mathcal{G}_{\varepsilon}^{j})$, for $j\in\{1,2\}$, the coupling function $(K^{j})^{-1}(\cdot,w)$ can be extended into a discrete meromorphic function $(K^{j}_{s})^{-1}(\cdot,w)$ on the graph $(\mathcal{G}_{\varepsilon}^{j})^{s}$ obtained by gluing together $\mathcal{G}_{\varepsilon}^{j}$ and $s(\mathcal{G}_{\varepsilon}^{j})$ (the reflection of $\mathcal{G}_{\varepsilon}^{j}$ along the flat part of $\partial U$). Then, $(K^{j}_{s})^{-1}(\cdot,w)$ is a discrete meromorphic function on $B((\mathcal{G}_{\varepsilon}^{j})^{s})$, with mixed Dirichlet/Neumann boundary conditions and two poles of respective residues $1/\pi$ and $-1/\pi$ at $w$ and $s(w)$, hence it coincides with the difference of the two coupling functions $H^{s}(\cdot,w)-H^{s}(\cdot,s(w)$ with the same mixed boundary conditions. Indeed, the difference

### v2

@@ Appendix B Inverse Kasteleyn matrix in Temperleyan domains
In Appendices B and C we present proofs of the behaviour of the inverse Kasteleyn Matrix in the small mesh limit - that it can be written as a sum of functions expressed in terms of the continuum Green function. These are technical proofs and are similar to computations in, for example, [14]. In this section we deal with ordinary Temperleyan domains, which is needed for our folded dimer model, and in particular show that the statement of the behaviour of the inverse Kasteleyn matrix extends to certain neighbourhoods of the boundary. In Appendix C we deal with piecewise Temperleyan domains, which is needed for our shifted dimer model.
In this section, the inverse Kasteleyn matrix on Temperleyan domains will often be called the coupling function as in [14]. We use the notation and setting of Section 3. We start by recalling some well-known facts on the coupling function and discrete holomorphic functions. The acquainted reader can jump directly to the statements of Lemma B.2 and Lemma B.3.
@@
where $G=\Delta^{-1}$ is the Green function. The function $G$ takes real values because the Laplacian $\Delta$ takes real values. Further, since the Laplacian has a block structure (it is non zero only from $B_{0}$ to $B_{0}$ and $B_{1}$ to $B_{1}$), $G$ has the same block structure. Hence, since $w\in W_{0}$, $w\pm\frac{\varepsilon}{2}\in B_{0}$ while $w\pm\frac{i\varepsilon}{2}\in B_{1}$, so $\tilde{K}^{-1}(b,w)$ is real for $b\in B_{0}$ and pure imaginary for $b\in B_{1}$ (justifying the terminology “real” and “imaginary” part).
Let us define $\partial^{out}B_{0}$ (the full red and blue circles on Figure 16) to be the vertices in $B_{0}(\varepsilon\mathbb{Z}^{2})\setminus B_{0}$ (where $B_{0}(\varepsilon\mathbb{Z}^{2})$ are the vertex of the dual graph of $\varepsilon\mathbb{Z}^{2}$ i.e. the faces of $\varepsilon\mathbb{Z}^{2}$) at distance $\varepsilon$ from a vertex of $B_{0}$ (see Kenyon, section 4.1) and $\partial^{out}B_{1}=\{b_{0}\}$ (full red square on Figure 16). Define $\overline{B_{0}}=B_{0}\cup\partial^{out}B_{0}$, $\overline{B_{1}}=B_{1}\cup\partial^{out}B_{1}$ and $\overline{B}=\overline{B_{0}}\cup\overline{B_{1}}$. If we extend $b\in B_{0}\to\tilde{K}^{-1}(b,w)$ by 0 on the vertices of $\partial^{out}B_{0}$, it is a harmonic function on $B_{0}$ except at two points: $\Delta\tilde{K}^{-1}(\cdot,w)_{|B_{0}}=\delta_{w+\varepsilon/2}-\delta_{w-\varepsilon/2}$. Indeed, for $b\in B_{0}$,
@@
as a function of $b\in B_{0}$. Observe that this already fully characterizes the real part: there is a unique function which satisfies (45) and is 0 on $\partial^{out}B_{0}$ (by the maximum principle). If we extend the imaginary part $b\in B_{1}\to\tilde{K}^{-1}(b,w)$ by 0 on $b_{0}$, the imaginary and real part are harmonically conjugated: $\tilde{K}\tilde{K}^{-1}=I$ implies that for $w^{\prime}\neq w$,
@@
In our case (and unlike Kenyon who considers multiply connected domains), the imaginary part is fully defined by saying that it is the harmonic conjugate of the real part and it is 0 at the removed vertex $b_{0}$: by discrete integration along paths which avoid the singularity at $w$, the values of $\tilde{K}^{-1}(\cdot,w)$ on $B_{1}$ can be uniquely recovered from the values on $B_{0}$ and the boundary conditions at $b_{0}$. We summarize all this in a lemma:
@@
  • its real part is the (unique) harmonic function on $\overline{B_{0}}$ which is 0 on $\partial^{out}B_{0}$ (Dirichlet boundary conditions) and satisfies
@@
We are now ready to give estimates of the coupling function. Let $\tilde{g}(u,v)$ be the analytic function of $v$ whose real part is the Dirichlet Green function $g(u,v)$ on $U$. Also define, as in Section 6.3 of [16],
@@
[…]
@@
We now show that the extended coupling function $b\in B^{s}\to\tilde{K}^{-1}_{s}(b,w)$ coincides with the difference of two coupling functions on $B^{s}$. Since $\mathcal{G}_{\varepsilon}^{s}$ is a Temperleyan approximation of $U\cup s(U)$ if we remove the vertex $b_{0}$, by Lemma B.1 for fixed $w\in W_{0}\cup s(W_{0})$, the coupling function $b\in B^{s}\to H^{s}(b,w)$ is uniquely defined: its real part has 0 boundary conditions on $\partial^{out}B_{0}\cup s(\partial^{out}B_{0})\cup\{b_{l},b_{r}\}\setminus\partial^{flat}B_{0}$ and satisfies $\Delta H^{s}(\cdot,w)=\delta_{w+\varepsilon/2}-\delta_{w-\varepsilon/2}$ at all inner points $b\in B^{s}_{0}$. Its imaginary part is 0 on $b_{0}$ and is harmonically conjugated with the real part except at $w$. We claim that for all $b\in B^{s}_{0}$,
@@
Observe that we did not impose any condition at $s(b_{0})$ for $H^{s}(b,w)-H^{s}(b,s(w))$, but since equality (50) holds it must be 0 a posteriori. This can be checked using symmetry arguments.
@@
If we denote by $\widetilde{g^{s}}(u,v)$ the analytic function of $v$ whose real part is the Dirichlet Green function $g^{s}(u,v)$ on $U\cup s(U)$, we define $F_{+}^{s}$ and $F^{s}_{-}$ similarly to the definition before Lemma B.2. If $w$ is at distance $O(\varepsilon)$ from $u$ and $b$ is at distance $O(\varepsilon)$ from $v$,
@@
where the $O(\varepsilon^{2})$ is uniform once $\delta$ is fixed. Observe now (this is the continuous equivalent of what we have just done, i.e. Schwarz reflection principle in the continuum) that $\widetilde{g^{s}}(u,v)-\widetilde{g^{s}}(s(u),v)=\tilde{g}(u,v)$ when $u,v\in U$, because the left-hand side is analytic in $v$ on $U$, its real part $g^{s}(u,v)-g^{s}(s(u),v)$ has the same pole as $g(u,v)$ in $v$ (by definition as a Green function) and it is 0 on the boundary of $U$. On the flat portion, it is true by symmetry: if $v$ is on the reflection axis, $g^{s}(u,v)=g^{s}(s(u),s(v))=g^{s}(s(u),v)$. On the rest of the boundary both terms of the difference vanish since $v$ is also on the boundary of $U\cup s(U)$. Hence, $F_{+}^{s}(u,v)-F_{+}^{s}(s(u),v)=F_{+}(u,v)$ and $F_{-}^{s}(u,v)-F_{-}^{s}(u,v)=F_{-}(u,v)$ which concludes the proof in the case $w\in W_{0}$; Lemma B.2 holds up to the boundary.
To deal with the case $w\in W_{1}$, the same argument works but now $\tilde{K}^{-1}(\cdot,w)_{|B_{0}}$ is the unique function on $B_{0}$ which is 0 on $\partial^{out}B_{0}$ and whose Laplacian is $i\delta_{b=w+i\varepsilon/2}+i\delta_{w-i\varepsilon/2}$ by equation (45) applied when $w\in W_{1}$ instead of $W\in W_{0}$. The remaining changes are straightforward. ∎
@@
In this final Appendix, we continue our analysis of the inverse Kasteleyn matrix in the small mesh limit, and in particular we deal with piecewise Temperleyan domains, which is needed for our shifted dimer model.
We use the setting of Section 3: $U^{r}$ is a simply connected open set symmetric with respect to the horizontal axis, $U=U^{r}\cap(\mathbb{R}\times\mathbb{R}_{>0})$. The boundary of $U$ can be partitioned in two connected arcs: $\partial U=D\sqcup N$ where $N=\partial U\cap(\mathbb{R}\times\{0\})$ is the intersection with the horizontal axis. Recall that we defined two graphs approximating $U$: the Temperleyan approximation $\mathcal{G}_{\varepsilon}^{1}$ with Dirichlet boundary conditions (and a vertex $b_{0}$ removed) and the piecewise Temperleyan approximation $\mathcal{G}_{\varepsilon}^{2}$ with two convex white corners $v_{1}^{*},v_{2}^{*}$. In both cases, we use the usual Kasteleyn phases of discrete holomorphy described in Equation (3), and denote by $K^{1},K^{2}$ the corresponding Kasteleyn matrices on $\mathcal{G}_{\varepsilon}^{1},\mathcal{G}_{\varepsilon}^{2}$. Both $\mathcal{G}_{\varepsilon}^{1}$ and $\mathcal{G}_{\varepsilon}^{2}$ have at least one dimer configuration (for example because from any loops and arcs configuration $\omega\in\Omega(\mathcal{G}_{\varepsilon})$ we can obtain two dimer covers of $\mathcal{G}_{\varepsilon}^{1}$ and $\mathcal{G}_{\varepsilon}^{2}$), hence $K^{1}$ and $K^{2}$ are invertible by [19]. Moreover, the asymptotics of $(K^{j})^{-1}$ on piecewise Temperleyan domains are computed and expressed in terms of $f_{i}^{j}(u,v)$ for $i\in\{0,1\},~j\in\{1,2\}$. These are the unique functions which are analytic in $u$ for fixed $v$, have a simple pole of residue $1/\pi$ at $u=v$ with no other pole on $\bar{U}$, are bounded near $v_{1}^{*},v_{2}^{*}$ (for the $j=2$ case), and have the following boundary conditions:
@@
where $\lambda=e^{i\frac{\pi}{4}}$ and $F^{\varepsilon}_{\mathbb{C},w}$ is the unique discrete holomorphic function on $\varepsilon\mathbb{Z}^{2}$ tending to 0 at infinity and with a pole $\lambda$ at $w$, which is known to be asymptotically equal to $\frac{\lambda}{2\pi(b-w)}$ when $\varepsilon\to 0$ and $|b-w|\geq\eta$ for some $\eta>0$ independent of $\varepsilon$. See Section 5 of [19] for details. The error term $\Delta$ is also present in the original statement of Lemmas B.2 and B.3, but we included it in the $o(\varepsilon)$ since we never use these lemmas for close points and for all $\eta>0$, for all $b,w$ such that $|b-w|\geq\eta$,
@@
As in Lemma B.3, when $U$ has some specific boundary conditions, the asymptotic estimates hold up to the boundary.
@@
The same argument as in the proof of Lemma B.3 can be applied: we skip most of the details. For fixed $w\in W(\mathcal{G}_{\varepsilon}^{j})$, for $j\in\{1,2\}$, the coupling function $(K^{j})^{-1}(\cdot,w)$ can be extended into a discrete meromorphic function $(K^{j}_{s})^{-1}(\cdot,w)$ on the graph $(\mathcal{G}_{\varepsilon}^{j})^{s}$ obtained by gluing together $\mathcal{G}_{\varepsilon}^{j}$ and $s(\mathcal{G}_{\varepsilon}^{j})$ (the reflection of $\mathcal{G}_{\varepsilon}^{j}$ along the flat part of $\partial U$). Then, $(K^{j}_{s})^{-1}(\cdot,w)$ is a discrete meromorphic function on $B((\mathcal{G}_{\varepsilon}^{j})^{s})$, with mixed Dirichlet/Neumann boundary conditions and two poles of respective residues $1/\pi$ and $1/\pi$ at $w$ and $s(w)$, hence it coincides with the difference of the two coupling functions $H^{s}(\cdot,w)-H^{s}(\cdot,s(w)$ with the same mixed boundary conditions. Indeed, the difference